OpenStudy (anonymous):

Essay; show all work. A community 5K run will award $50 to the winner. 45 people enter the race, and they each pay an entry fee of $20. Assuming they are all equally likely to win, what is a fair price for the competition? Round to the nearest cent

OpenStudy (anonymous):

\(\frac{50}{45}\)

OpenStudy (anonymous):

if each person pays \(\frac{50}{45}\) then they will raise \(\frac{50}{45}\times 45=50\) dollars

OpenStudy (anonymous):

It says they will be awarded 50 dollars and pay 20 dollars to get in. Im confused on how 50 dollars would be right